# W-RAG: Weakly Supervised Dense Retrieval in RAG for Open-domain Question Answering

### v1

@@ 4.4. Evaluation
We use Recall as our primary metric to evaluate the retrieval performance, as it focuses on whether the relevant passage is included within a specific ranking range. We also report MRR (Mean Reciprocal Rank), which is similar to Recall but also considers the position at which the relevant passage is ranked.
To gauge the alignment between machine-generated answers and ground-truth answers, we use F1, Rouge-L (Lin, 2004), and BLEU-1 (Papineni et al., 2002). These metrics collectively measure the overlap and sequence matching between generated and actual answers. We evaluate OpenQA performance on our uniformly random sampled test set, which contains 2,000 questions for each datasets. For questions with multiple correct answers in the dataset, we select the answer with the highest cumulative score across all OpenQA metrics.
Retriever MSMARCO QnA NQ SQuAD WebQ
@@
ColBERTinit 0.0163 0.0120 0.0220 0.0866 0.0093 0.0224 0.0504 0.0063 0.0158 0.0031 0.0001 0.0004
BM25 0.2781 0.1647 0.4580 0.5686 0.0869 0.3575 0.2978 0.0400 0.2011 0.2737 0.0285 0.1336
Contriever 0.2826 0.1585 0.4901 0.5524 0.0988 0.3585 0.2221 0.0294 0.1606 0.2990 0.0165 0.1304
ReContriever 0.2908 0.1589 0.5093 0.5831 0.1118 0.3952 0.2150 0.0284 0.1597 0.3022 0.0194 0.1370
Trained on Groundtruth Data
$\text{DPR}_{base}$ 0.3237 0.2117 0.4994 0.5608 0.1174 0.3619 0.1786 0.0199 0.1054 0.4028 0.0561 0.2059
ColBERTbase 0.3484 0.2097 0.5112 0.5925 0.1316 0.3082 0.2310 0.0469 0.1072 0.1157 0.0180 0.0501
DPRReCon 0.3652 0.2284 0.5913 0.6632 0.1516 0.4715 0.2829 0.0393 0.2072 0.4155 0.0640 0.2262
Trained on W-RAG Data
$\text{DPR}_{base}$ 0.2981 0.1809 0.4813 0.5302 0.1012 0.3364 0.1781 0.0210 0.1020 0.3377 0.0478 0.1656
ColBERTbase 0.3116 0.1973 0.4904 0.5229 0.0982 0.2672 0.2265 0.0445 0.1029 0.0831 0.0153 0.0326
DPRReCon 0.3496 0.2023 0.5887 0.6301 0.1311 0.4267 0.2811 0.0382 0.2001 0.3610 0.0499 0.1880
Table 2. Retrieval Performance after training with W-RAG or ground-truth data. : Performance improvements are statistically significant with a ttest of $p<0.05$.
Retriever MSMARCO QnA NQ SQuAD WebQ
@@
BM25 0.1694 0.4464 0.7577 0.088 0.353 0.7941 0.0449 0.2069 0.6461 0.0217 0.1155 0.5765
Llama3-8B 0.5239 0.6614 0.7517 0.1542 0.4224 0.7303 0.0805 0.3088 0.6774 0.0494 0.1887 0.5502
Table 3. BM25 retrieves the top 100 passages, which are then reranked by Llama3-8B. The reranked lists are used as weak labels for training a dense retriever.
@@ 5. Experimental Results
In this section we discuss the final OpenQA performance, retrieval performance of trained retrievers, the quality of weak labels, and ablation studies.
@@ 5.1. Main Results
We do not claim that W-RAG outperforms state-of-the-art RAG methods or achieves the best retrieval performance on any of the four datasets. Instead, we highlight the performance improvements of weakly trained retrievers compared to baselines, and demonstrate that the gap between weakly supervised and ground-truth trained retrievers is relatively small. This relationship is evident in both the retrieval performance and the final RAG for OpenQA performance.
The goal of W-RAG is to enhance LLM generation quality by providing relevant evidence passages. Tabel 1 presents the main OpenQA results using different retrievers in the RAG pipeline, evaluated on 2,000 test questions for each dataset. For these experiments, only the top 1 retrieved passage is added to the prompt, and the LLM is limited to generating 20 tokens. The Naive and Groundtruth baselines represent the lower and upper bounds, respectively. The Naive baseline excludes any supplementary passage, while Groundtruth includes the best passage. Any retriever that retrieves relevant passages should outperform the Naive baseline, while surpassing the Groundtruth baseline would be very unexpected. Results show consistent improvements with W-RAG trained retrievers across all datasets, with most of them being statistically significant. The relatively small gap between W-RAG trained and ground-truth trained retrievers suggests that W-RAG data approaches the quality of human-labeled data.
One baseline, ColBERTinit (initialized with bert-base-uncased), underperforms the Naive baseline, likely due to irrelevant passages being retrieved, introducing noise that misguides Llama3. We notice that for NQ, SQuAD, and WebQ, all the OpenQA metrics are small. This is due to the brevity of ground-truth answers (average length  2 words). For example, in response to the NQ question ”When did Big air Snowboarding become an Olympic sport?”, Llama3 outputs the correct answer: ”Big Air Snowboarding was first added to the 2018 Winter Olympics.” Since the ground-truth answer given by NQ is just ”2018”, standard OpenQA metrics like F1 and Rouge-L would yield low scores due to the denominator of the Precision term being the length of the generated answer, while the numerator is n-gram overlap which is at most 1 in this case. Despite this, these metrics remain valuable; if Llama3’s answer omitted ”2018,” the score would be zero. The absolute metric values are less important than the relative differences across settings. Although token length could be treated as a hyperparameter, we limited generation to 20 tokens across datasets for consistency.
@@ 5.2. Retrieval Results
The retrieval results of different trained retrievers are presented in Table 2. These results are based on 2,000 test questions for each trained or baseline retriever. On the MSMARCO QnA dataset, all retrievers trained on W-RAG data outperformed all baselines, while DPRReCon consistently outperformed all unsupervised baselines, except for SQuAD where BM25 performed best. As expected, DPR trained on ReContriever showed better retrieval performance than DPR trained from scratch. Similar to the OpenQA task, W-RAG trained retrievers slightly underperform compared to those trained on ground-truth data.
Although there appears to be a positive correlation between retrieval performance and OpenQA performance, this relationship is noisy and sometimes inconsistent. For instance, W-RAG trained ColBERTbase achieved nearly 4% higher Recall@1 than ReContriever on MSMARCO QnA, ranking the relevant passage at top for about 80 more questions. However, when these top 1 passages are used for OpenQA, ColBERTbase performed slighly worse than ReContriever. Similarly, ground-truth trained DPRReCon significantly outperformed W-RAG trained DPRReCon in Recall@1 on MSMARCO QnA, yet their OpenQA performance was almost identical when using the top-ranked passage.
These findings suggest that the traditional definition of relevance may not directly impact the answer quality in RAG systems. W-RAG’s top passage is selected based on its ability to elicit the correct answer from the LLM, while ground-truth passages are typically chosen based on term overlap, semantic similarity, or human judgment. Contrary to expectations, better retrieval does not always lead to better OpenQA. This raises the question: “Is the relevance definition in existing datasets truly effective for evaluating retrievers in RAG systems?”
@@ 5.3. W-RAG Labels
After generating all of W-RAG data, we examine its retrieval performance and the quality of the weakly labeled rank lists. In Table 3, a significant gap between BM25 and Llama3’s reranking performance at Recall@1 is evident, which justifies our choise of only using the top 1 passage as the positive when training different retrievers. We also observe that Llama3’s performance declines as the number of passages increases. This is likely due to the fact that starting from around the 50th position, the difference in answer likelihood between adjacent passages diminishes to negligible values, such as $\exp(-3.24815)-\exp(-3.24899)\approx 3\times 10^{-5}$.
Prompt R@1 R@5 F1 Rouge-L BLEU
BM25 0.1343 0.4128 0.2797 0.2436 0.0640
UPR 0.1970 0.4970 0.2898 0.2538 0.0677
Zero-shot 0.5027 0.6289 0.4300 0.3924 0.1876
One-shot 0.5187 0.6690 0.4358 0.3976 0.1939
Two-shot 0.5137 0.6840 0.4256 0.3858 0.1812
Table 4. Llama3-8B-Instruct’s reranking performance using different prompts on 500 validation questions from the MSMARCO QnA dataset. Answers are generated using only the top 1 reranked passage.
@@ 5.4. Ablation Study
We assessed the quality of W-RAG data generated by different LLMs, specifically their ability to rerank passages based on answer likelihood. This is illustrated in Figure 3, showing the reranking performance of various LLMs on 500 questions from the validation set of MSMARCO QnA. Although the LLMs notably differs at Recall@5, they exhibit consistent trends with each other. We notice Llama3-8B-Instruct is weak when compared to other examined LLMs. However, since their recall distributions are similar, any of these LLMs can serve as the reranker. We chose to use Llama3-8B-Instruct for all our experiments.
# F1 Rouge-L BLEU-1 Latency (sec/infer)
1 0.3397 0.3074 0.2559 1.043
3 0.3744 0.3395 0.2818 1.270
5 0.3767 0.3427 0.2835 1.454
10 0.3864 0.3524 0.2911 2.134
Table 5. Adding different # of evidence passages to the prompt when doing OpenQA using DPRReCon on 2,000 test questions from MSMARCO QnA.
Using Llama3-8B-Instruct, we evaluated reranking and OpenQA performance, as presented in Table 4. UPR (Sachan et al., 2022a), which ranks passages based on question likelihood, serve as an unsupervised approach for generating training labels. With weak supervision of ground-truth answers, W-RAG’s zero-shot, one-shot, and two-shot prompts significantly outperforms BM25 and UPR in both reranking and OpenQA metrics. However, the differences between the various shot prompts are not very consistent and are not significant. We hypothesize that this inconsistency may have stemmed from the LLM misinterpreting example passages as the relevant passage, or if the question types are drastically different between the example and the given question, thus introducing noise and reducing result reliability. Due to this, we conducted all subsequent experiments using zero-shot prompts, shown in Figure 2.
During the final OpenQA, we also study the impact of giving more evidence passages, shown in Table 5. We observe a consistent and steady increase in all OpenQA metrics when more evidence passages are inserted into the prompt. This behavior is expected because the more passages supplied, the better chance a good passage is within them. However, latency also grows as we increase the number of passages.
@@ 6. Conclusions and Future Work
This paper introduces a general framework, W-RAG, for extracting weak labels from question-answer pairs to address the scarcity of training data for dense retrieval in RAG for OpenQA. W-RAG reranks the top-$K$ initially retrieved passages by the probability of LLMs generating the OpenQA answer conditioned on each question-passage pair. This ranking score computation method aligns well with the latest study (Cuconasu et al., 2024), which states that the retrieved passage should answer the question; otherwise, performance will be negatively impacted. Extensive experimental results on four public OpenQA datasets demonstrate the effectiveness of W-RAG.
In future work, we plan to explore which types of passages most effectively enhance RAG’s performance in OpenQA, as indicated by (Cuconasu et al., 2024), where even randomly sampled tokens were beneficial in some cases. Understanding the types of passages preferred by LLMs will allow us to design more effective dense retrieval methods, including new structures and evaluation metrics. Additionally, the compression of retrieved passages warrants further study, as directly feeding all retrieved passages to LLMs not only increases computational complexity but also introduces significant noise. Finally, even with a ground-truth evidence passage, RAG can still produce incorrect answers, known as hallucinations; enhancing the robustness of RAG in OpenQA is another promising direction for future research.

### v2

@@ 4.4. Evaluation
To evaluate retrieval performance, we use Recall as our primary metric, as it reflects whether the relevant passage is included within a specific top-$k$ ranking range. We also report Mean Reciprocal Rank (MRR), which measures the inverse of the rank at which the first relevant passage appears. Unlike Recall, which only considers whether a relevant passage is retrieved, MRR also takes into account its position in the ranked list.
For OpenQA metrics, we use F1, Exact Match (EM), and BLEU-1 (Papineni et al., 2002) to assess the alignment between LLM-generated answers and ground-truth answers. F1 measures the token-level overlap between generated and reference answers. EM evaluates whether the generated answer exactly matches any of the ground-truth answers. BLEU-1 computes unigram precision, measuring the proportion of overlapping words between the LLM-generated answer and the reference answers without considering word order or recall. We evaluate OpenQA performance on our randomly sampled test set, which contains 2,000 questions per dataset. For questions with multiple ground-truth answers, the generated answer is compared with all references, and the highest resulting score is used for evaluation.
Retriever NQ SQuAD WebQ MSMARCO QnA
@@
Without fine-tuning
BM25 56.86 8.69 35.75 29.78 4.00 20.11 27.37 2.85 13.36 27.81 16.47 45.80
Contriever 55.24 9.88 35.85 22.21 2.94 16.06 29.90 1.65 13.04 28.26 15.85 49.01
ColBERTinit 8.66 0.93 2.24 5.04 0.63 1.58 0.31 0.01 0.04 1.63 1.20 2.20
ReContriever 58.31 11.18 39.52 21.50 2.84 15.97 30.22 1.94 13.70 29.08 15.89 50.93
Fine-tuned with ground-truth data
DPRbase 56.08 11.74 36.19 17.86 1.99 10.54 40.28 5.61 20.59 32.37 21.17 49.94
ColBERTbase 59.25 13.16 30.82 23.10 4.69 10.72 11.57 1.80 5.01 34.84 20.97 51.12
DPRReCon 66.32 15.16 47.15 28.29 3.93 20.72 41.55 6.40 22.62 36.52 22.84 59.13
Fine-tuned with W-RAG data
DPRbase 53.02 10.12 33.64 17.81 2.10 10.20 33.77 4.78 16.56 29.81 18.09 48.13
ColBERTbase 52.29 9.82 26.72 22.65 4.45 10.29 8.31 1.53 3.26 31.16 19.73 49.04
DPRReCon 63.01 13.11 42.67 28.11 3.82 20.01 36.10 4.99 18.80 34.96 20.23 58.87
Table 2. Retrieval Performance after training with W-RAG or human-labeled data. Bold indicates the best performance; Underlined indicates the second best. : Performance improvements are statistically significant with a t-test of $p<0.05$.
Retriever NQ SQuAD WebQ MSMARCO QnA
@@
BM25 8.8 35.3 79.41 4.49 20.69 64.61 2.17 11.55 57.65 16.94 44.64 75.77
Llama3-8B-Instruct 15.42 42.24 73.03 8.05 30.88 67.74 4.94 18.87 55.02 52.39 66.14 75.17
Table 3. BM25 retrieves the top 100 passages, which are then reranked by Llama3-8B-Instruct. The reranked lists are used as weak labels for training a dense retriever.
@@ 5. Experimental Results
In this section, we discuss the final OpenQA performance, the retrieval performance of fine-tuned retrievers, the quality of weak labels, and the results of ablation studies.
@@ 5.1. Main Results
The goal of W-RAG is to draw weak training signals from the downstream task (such as OpenQA) of an LLM, and fine-tune the retriever to prioritize passages that most benefit the task. In other words, W-RAG aims to improve OpenQA performance by aligning retrieval with answer generation. We highlight the performance improvements of weakly supervised retrievers compared to baselines, and demonstrate that the gap between weakly supervised and ground-truth supervised retrievers is relatively small. This trend is observed in both retrieval metrics and final performance in the RAG-based OpenQA setting. Table 1 presents the main OpenQA results using different retrievers in the RAG pipeline, evaluated on 2,000 test questions for each dataset. For these experiments, the top 5 retrieved passages are added to the prompt, as shown in Figure 3. The “Naive” and “Ground truth” represent the lower and upper bounds of OpenQA performance under our setup, respectively. The “Naive” baseline excludes any supplementary passage, while “Ground truth” includes the ground-truth passage. Any retriever that retrieves relevant passages should outperform the “Naive” baseline, while surpassing the “Ground truth” upper bound would be very unexpected. Figure 4 provides a more intuitive comparison of the OpenQA results. Across all datasets, W-RAG-tuned retrievers consistently outperform baselines within both the DPR and ColBERT model families. Most of these improvements are statistically significant. Furthermore, the relatively small performance gap between W-RAG-tuned and ground-truth-tuned retrievers suggests that the weak labels generated by W-RAG approach the quality of human-labeled data.
One baseline, ColBERTinit (initialized with bert-base-uncased), underperforms the “Naive” baseline, likely due to irrelevant passages being retrieved, introducing noise that misguides the LLM. For MSMARCO QnA, the EM values are relatively smaller because the ground-truth answers are generally much longer than NQ, SQuAD, or WebQ, and are therefore more challenging for the LLM to generate exactly the same words as the ground-truth answer.
@@ 5.2. Retrieval Results
The retrieval results of different fine-tuned retrievers are presented in Table 2. These results are based on 2,000 test questions for each fine-tuned or baseline retriever. On NQ, WebQ, and MSMARCO QnA datasets, all retrievers fine-tuned on W-RAG data outperformed their corresponding baselines, while DPRReCon consistently outperformed all unsupervised baselines, except for SQuAD where BM25 performed best. As expected, DPR fine-tuned from ReContriever showed better retrieval performance than DPRbase, which was fine-tuned from the base BERT model. Similar to the OpenQA task, retrievers fine-tuned with W-RAG data slightly underperform those fine-tuned with ground-truth data.
Although there is a general positive correlation between retrieval performance and OpenQA performance, the relationship is noisy and sometimes inconsistent. For instance, on SQuAD, BM25 achieved a much higher Recall@5 (20.11) compared to ColBERTbase fine-tuned with W-RAG data (10.29), meaning BM25 ranked the relevant passage among the top 5 for approximately 200 more questions. However, in OpenQA, using the top 5 passages from ColBERTbase led to better answer accuracy than using those from BM25. Similarly in NQ, DPRReCon fine-tuned on W-RAG achieved significantly higher retrieval performance than ColBERTbase (Recall@5 of 42.67 vs. 26.72), their OpenQA performances were relatively close.
This discrepancy may be partly due to the prompt instructing Llama3.1-8B-Instruct to fall back on its internal knowledge when the retrieved passage cannot answer the question. Nevertheless, these findings suggest that traditional notions of relevance may not directly translate into improved quality of answers in RAG systems. Retrievers trained on W-RAG data select top passages based on their ability to elicit the correct answer from the LLM, while ground-truth passages are typically chosen based on term overlap, semantic similarity, and human judgment. Contrary to expectations, stronger retrieval metrics do not always lead to better OpenQA results. This raises a critical question: Is the relevance definition in existing datasets truly effective for evaluating retrievers in RAG systems?
@@ 5.3. W-RAG Labels
After generating the W-RAG data, we evaluate its retrieval performance and assess the quality of the weakly labeled ranking lists. As shown in Table 3, there is a significant gap in Recall@1 between BM25 and the reranking performance of Llama3-8B-Instruct. This observation supports our decision to use only the top-1 passage as the positive example when fine-tuning different retrievers. We also observe that the quality of the reranked list deteriorates at higher recall values. This decline occurs because the difference in answer likelihood between adjacent passages becomes negligible beyond a certain rank. For example, the difference in answer likelihood between the 50th and 51st ranked passages is approximately $\exp(-3.24815)-\exp(-3.24899)\approx 3\times 10^{-5}$, which is very insignificant. For passages that do not meaningfully increase the LLM’s likelihood of generating the correct answer, the assigned likelihood scores remain similarly low across the ranking.
@@ 5.4. Ablation Study
To study the quality of W-RAG data as generated by different LLMs, we experimented with several LLMs (Jiang et al., 2023a; Rivière and et al., 2024), as shown in Figure 5. The figure reports reranking performance on 500 questions randomly selected from the validation set of MSMARCO QnA. Although the LLMs differ slightly in Recall@5, they exhibit consistent trends. This consistency gives us confidence that the W-RAG method generalizes well to models at the 8B parameter scale. We leave the more extensive study with larger models for future work. We notice Llama3-8B-Instruct is relatively weaker when compared to other examined LLMs. Nevertheless, the recall distributions across LLMs are similar, and we conclude that any of these models can effectively serve as the agent for generating weak supervision signals. In the 8B parameter scale, this choice does not substantially impact the outcome of the trained retrievers and the downstream OpenQA performance.
We also examine the trade-off between OpenQA accuracy and latency on the NQ dataset by varying the number of top passages retrieved by DPRReCon, as shown in Table 4. We observe a consistent and steady improvement in all OpenQA metrics as more evidence passages are included in the prompt. This trend begins to plateau between 10 and 20 passages under our experimental setting. This behavior is expected: supplying more passages increases the likelihood that at least one of them contains the correct answer. However, latency increases linearly with the number of inserted passages, making it a trade-off between answer quality and inference speed.
# Passages F1 EM BLEU-1 Avg Latency (ms)
1 54.64 40.95 50.50 264.90
3 57.74 42.90 53.14 342.01
5 60.21 45.35 55.62 413.96
10 61.32 46.10 56.59 602.33
20 61.39 45.40 56.15 1005.50
Table 4. OpenQA performance on NQ using different numbers of top passages retrieved by DPRReCon fine-tuned with W-RAG data.
@@ 6. Conclusions and Future Work
This paper introduces a general framework, W-RAG, for extracting weak signals from question-answer pairs using an LLM to address the scarcity of training data for dense retrieval in RAG for OpenQA. W-RAG reranks the top-$k$ initially retrieved passages by the probability of LLMs generating the ground-truth answer conditioned on each question-passage pair. This ranking score computation method aligns well with a recent study (Cuconasu et al., 2024), which states that the retrieved passage should answer the question; otherwise, performance will be negatively impacted. Extensive experimental results on four public OpenQA datasets demonstrate the effectiveness of W-RAG. For future work, we plan to scale the generation of W-RAG data to a quantity comparable to that used in training state-of-the-art retrievers, and evaluate whether retrievers trained with W-RAG can surpass them in final OpenQA performance when integrated into a RAG system. As for future research directions, we plan to explore which types of passages most effectively enhance RAG’s performance in OpenQA, as indicated by (Cuconasu et al., 2024), where even randomly sampled tokens were beneficial in some cases. Understanding the types of passages preferred by LLMs will allow us to design more effective dense retrieval methods, including new structures and evaluation metrics. Additionally, the compression of retrieved passages warrants further study, as directly feeding all retrieved passages to LLMs not only increases computational complexity but also introduces significant noise. Finally, even with a ground-truth evidence passage, RAG can still produce incorrect answers, known as hallucinations; enhancing the robustness of RAG in OpenQA is another promising direction for future research.